RC meeting and Run Plan, Monday September 26th

From SBS wiki
Jump to: navigation, search


RC Meeting Information

Run Coordinator: Donald Jones

Location: Counting House Conference Room

Time: 3:00PM

Commissioning run plans

  • Beam line (1-2 shifts) Sept 30.
    • Beam centering and target alignment
    • Raster calibration
    • BCM and BPM calibration
  • Corrector adjustment procedure
  • GEMs run plan
  • Charge feedback system
  • Optics and calibration
  • Moller polarimeter (1 shift) Oct 7
  • Target commissioning
  • Bigbite commissioning (preshower, shower, GRINCH)
  • HCal commissioning

Subsystem status reports

  • Hall A Status -- Jessie Butler
  • Target status/timeline (training?) -- Arun
  • Beam line:
    • Monitors (calibration) -- Simona
    • Correctors -- Bogdan, Iuliia
  • Tentative run plan document -- Bogdan
  • Alarms
  • Analyzer -- Andrew
  • DAQ -- Alex
  • HCal -- Jiwan
  • Spectrometer magnets -- Jack
  • GEMs/Tracking -- Holly
    • Are new HV controls, alarms, gas flow monitoring procedures, EPCIS tracking in place?
    • Shift checklist ok?
  • Shower Calorimeter -- Provakar
  • Beam polarization -- Don


  • RC reported that the SBS and BB magnets were successfully ramped to 2100A and 750A respectively. Jack S. is working to increase lcw flow to SBS. He measured 50G in target chamber at high field and 5G at half field. Bogdan speculates that this non-linearity is due to being near the beam window.
  • Arun reported that a fourth oven heater is being installed and they will test the oven to see if they can get to 255 degrees C once it is complete. If the oven reaches 255, then the target ladder will be installed and the Ukraine and the reference cell and foil ladder will be installed in preparation for survey and alignment either Tuesday or Wednesday.
  • Simona showed raster uniformity study results. BPMs are not fast enough to resolve the carbon hole. She has iterated with Yves and finalized the plan to bring beam into the hall. Target experts will be on call but not required present.
  • Bogdan advocates for alternative plan with no target in case the target is not done in time.
  • Todd reported that the reference cell is ready for installation in the target ladder.
  • RC asks for detailed run plans from experts.